Differences between revisions 11 and 12
Revision 11 as of 2022-06-17 14:26:25
Size: 2996
Revision 12 as of 2022-06-17 14:31:37
Size: 3025
Deletions are marked like this. Additions are marked like this.
Line 62: Line 62:
 * Type-B: FC to SR3/SRM. Investigation on the mirror replacement in SRM.  * Type-B: FC to SR3/SRM. Geophone calibration in SR2. Investigation on the mirror replacement in SRM.


Zoom Meeting

CRY+VIS Interface Meeting

Date: 2022/6/17 13:00-13:20

Participants: Takahashi, Sato, Ikeda, Aso, Uchiiyama, Akutsu, Kimura, Hirata, Yokozawa

Post O3 Upgrade portal


  • IX: Heater control of the laser for OpLev.

  • EX:
  • IY: Leak test. Installation of the Windshield. Cooling of the ducts is ready.
  • EY: Vent in the next week.
  • Type-B: Stand-alone SM driver for SR3. FC to SR3/SRM in the next week. Investigation of the mirror replacement in SRM in the next week.
  • Type-Bp: Recovery of the alignment after the FC in PR3. Installation of the traverser limit SW in PR2. Installation of the MSB in PR3. Health check in the next week.
  • Type-C:


  • Replacement of flip cables: EXT and BF temp.
  • TM height setting: Following the cooling.
  • Oplev: Investigating the relation between laser and temperature. Adjustment for Type-Bp is necessary.
  • ACC:
  • Commissioning in air:
  • Commissioning in vacuum:
  • Acceptance check:
  • PF actuator: At the in-vacuum commissioning.
  • New vacuum pumps for V-chambers: Test of the IYV pump.
  • Heater control: Pending.
  • TMS:


VIS Meeting

Date: 2022/6/17 14:00-14:25

Participants: Takahashi, Uchiyama, Fabian, Tanaka, Hirata, Aso, Sato

Summary Report



  • Assembled the stand-alone SM driver. Connected it to the SF in SR3. It successfully worked.
  • SR2 new geophone: I noticed that the TF of IP-Y with the new geophone, had a depression around 1 Hz, so I calculated corrections to the geophone parameters. See klog.


  • In PR3, we removed the First Contact on Monday ad released the suspension. As expected, by locking it for cleaning, we disturbed the pitch of the mirror. However, the misalignment this time was different from others before (see a klog from today).
  • Ikeda-san and Hirata-san did in-air cabling for the Traverser.
  • In PR2 we installed the Traverser limit switches and new cable clamps. A few cables broke in the process, but they were fixed. We also did in-vacuum and in-air cabling for the Traverser.



  • The shorted part of the broken geophone pre-amplifier was a tantalum capacitor of 100uF on the positive power line.



  • Type-A: Monitoring during the vent in EY.
  • Type-B: FC to SR3/SRM. Geophone calibration in SR2. Investigation on the mirror replacement in SRM.
  • Type-Bp: Test of the traverser limit SW in PR2. Health check.
  • Type-C:



Travel Plans

(Date in Kamioka)

The week from 6/20

  • Takahashi:
  • Sato:
  • Hirata: 6/20-24

The week from 6/27

  • Takahashi: 6/27-7/1
  • Sato: 6/28-7/1
  • Hirata: 6/28-7/1

Next meeting

  • 2022/6/24(FRI)

KAGRA/Subgroups/VIS/VISMinutes20220617 (last edited 2022-06-17 14:31:37 by RyutaroTakahashi)